OpenStudy (anonymous):

@zordoloom 7^t=8^t+1

OpenStudy (anonymous):

Ok, start by taking the ln or log of both sides.

OpenStudy (anonymous):

that seems to be the part i keep getting stuck on i end up with tlog7=tlog8+log8

OpenStudy (anonymous):

ok, Let me write out the steps then.

OpenStudy (anonymous):

Okay thanks

OpenStudy (anonymous):

Actually know what. let me take a different approach.

OpenStudy (anonymous):

Divide both sides by 7^t

OpenStudy (anonymous):

So, 1=(8^(t+1))/(7^t)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Now if you rewrite (8^(t+1))/(7^t), rewrite it as 7^-t(8^(t+1)=1

OpenStudy (anonymous):

Are you following so far?

OpenStudy (anonymous):

Um yeah i think so

OpenStudy (anonymous):

Where at?

OpenStudy (anonymous):

Are you confused.

OpenStudy (anonymous):

So the t is now negative on the 7 because we're dividing right?

OpenStudy (anonymous):

Its negative because it moved from denominator to numerator.

OpenStudy (anonymous):

Okay i gotcha now

OpenStudy (anonymous):

Now, you want a common base. So raise both sides by e and take the log of both sides. Let me show you what i mean.

OpenStudy (anonymous):

e^((log(8)(t+1)-tlog(7))=1

OpenStudy (anonymous):

now take the ln of both sides.

OpenStudy (anonymous):

Now combine in terms of t t(ln(8)-ln(7))+ln(8)=0

OpenStudy (anonymous):

so, t=(-ln8)/(ln8-ln7))

OpenStudy (anonymous):

Okay let me see if i get the correct answer

OpenStudy (anonymous):

ok

OpenStudy (anonymous):



OpenStudy (anonymous):

Okay so when i punched it in i got -15.57 is that correct or does it just remain t=(-ln8)/(ln8-ln7))

OpenStudy (anonymous):

Both would be correct.

OpenStudy (anonymous):

Okay great! thanks again...

OpenStudy (anonymous):

If you want exact form, use the fraction looking example, if you want approximate, use the decimal.

OpenStudy (anonymous):

I'll probably end up using the exact form just in case

OpenStudy (anonymous):

Thats a good choice. Thats what i would leave it in.

OpenStudy (anonymous):

Best response?